Part 1:

Find the marginal utility of an additional unit of good x for a consumer with utility function u (z, y) x-y- when their current consumption level is 2 and = 16

Part 2:

Find the marginal rate of substitution between goods LaTeX: x and LaTeX: y for a consumer with utility function u (z, y) x-y- when their current consumption level is 2 and = 16
Answer #1

MUx = 2/3 (x) -1/3 y1/3

x=2 and y=16

MUx = 2/3 (2)-1/3 (16)1/3

= 2/3 [16/2]1/3

= 2/3 [8]1/3

=2/3 (2) = 4/3 (Marginal utility of good x)

MUy = 1/3 (x)2/3 (y)-2/3

= 1/3 (2)2/3(16)-2/3

= 1/3 (2/16)2/3

= 1/3 [1/8]2/3

= 1/3 (1/2)2

= 1/3 (1/4) = 1/12 (Marginal utility of good y)

Marginal rate of substitution =MUx/ MUy

= (4/3) / (1/12)

= (4)(4)= 16